# bibliography.bib

### v1

@@ 5.2 Elliptic PDE with Lognormal Random Coefficients – Direct Solver
In the last example we show how further memory gains can be obtained using quarter precision on the coarsest discretisation level. In this experiment, we again solve the problem (27) with $f\equiv 1$ on $D$ and the parameters in the coefficient function chosen as $s=1$ and $\sigma=1$. Unlike in the previous experiments, the initial discretisation parameter is chosen as $h_{0}=1/4$. In the adaptive MPML algorithm (Algorithm 4.1) we choose $k_{p}=0.4$. Iterative refinement with the Cholesky solver is used with the following settings: $qhhh$ on level 00, and $hhss$ on levels $2$ and $3$. This means that on level 00 corresponding to $h_{0}=1/4$, quarter precision (q43, see Section 2.1) is used for the Cholesky factorisation. To simulate quarter precision, the pychop package of Xinye Chen was used; see [pychop]. Figures 7 and 8 show the MPML MSE and simulated memory gain, respectively. They have been produced analogously to Figures 5 and 6, respectively. We observe that for the MSE tolerance $2\times 10^{-6}$, the MPML relative error is $11\%$ bigger compared to standard MLMC with the same number of samples while achieving a total memory gain of $\approx 3.5$. Recall that the cost gain is simulated by counting non-zeros computed and stored in the process of solving the linear system $Ax=b$.
@@
We have not encountered overflow in any of the examples presented in this section. It is important to note that due to the fact that the coefficients of the PDE are lognormally distributed, overflow can occur with low probability. In the case where overflow occurs, scaling or shifting techniques can be used; see [higham2019squeezing].
@@ 6 Conclusion
Multilevel sampling methods have proven to be powerful tools for uncertainty quantification, offering significant performance improvements by efficiently redistributing computational work across a hierarchy of models. In this paper, we demonstrated how leveraging computational error in coarser levels can further enhance the efficiency of these methods in high-performance computing applications. As a use case, we have developed an adaptive algorithm to determine the minimum required computational accuracy for each level in the multilevel Monte Carlo method.
Through two practical examples, we showcased the potential of our approach to obtain significant cost gains. Using a low-precision sparse direct solver with iterative refinement, we achieved a simulated memory gain up to 3.5×3.5$\times$3.5 ×, while employing a MINRES iterative solver yielded a speedup of 2×2$\times$2 × in floating point operations. These results highlight the significant potential for energy-aware scientific computing. The potential for future work lies in generalising this approach to other uncertainty quantification frameworks, such as multilevel Markov chain Monte Carlo or the multilevel stochastic collocation method and exploring its broader applications in scientific computing. Efficient parallel implementation of low-precision solvers within multilevel sampling method is of great interest and promises to bring further speedup.

### v2

@@ 5.2 Elliptic PDE with lognormal coefficients – direct solver
In the last example, we show how further memory gains can be obtained using quarter precision on the coarsest level. For this experiment, we again solve (27) with $f\equiv 1$ on $D$, $s=1$ and $\sigma=1$, but now we choose the coarsest mesh size to be $h_{0}=1/4$. To determine the effective precision $\delta_{l}$ in (24) we choose $k_{p}=0.4$. Iterative refinement with the Cholesky solver is used with the following settings: $qhhh$ on level 00, and $hhss$ on levels $1$ to $3$. This means that on level 00 where $h_{0}=1/4$, quarter precision (q43, see Section 2.1) is used for the Cholesky factorisation. To simulate quarter precision, the pychop package of Xinye Chen was used; see [pychop]. Figures 7 and 8 show the MSE for MPML and MLMC and the simulated memory gain, respectively. They have been produced analogously to Figures 5 and 6. We observe that for the MSE tolerance $2\times 10^{-6}$, the MPML relative error is $11\%$ bigger compared to standard MLMC error with the same number of samples, while we are able to achieve a total memory gain of $\approx 3.5$. Recall that the cost gain is simulated by counting non-zeros computed and stored in the process of solving the linear system $Ax=b$.
@@
We have not encountered overflow in any of the examples presented in this section when working with the lower precisions. It is important to note that due to the fact that the coefficients of the PDE are lognormally distributed, overflow can occur with low probability. In the case where overflow occurs, scaling or shifting techniques can be used; see [higham2019squeezing].
@@ 6 Conclusion
Multilevel sampling methods have proven to be powerful tools for uncertainty quantification, offering significant performance improvements by efficiently redistributing computational work across a hierarchy of models. In this paper, we demonstrated how leveraging computations of lower accuracy on coarser levels can further enhance the efficiency of these methods in high-performance computing applications. As a use case, we have developed an adaptive algorithm to determine the minimum required computational accuracy for each level in the multilevel Monte Carlo method.
Through two practical examples, we showcased the potential of our approach to obtain significant cost gains. Using a low-precision sparse direct solver with iterative refinement, we achieved a simulated memory gain of up to 3.5×3.5$\times$3.5 ×, while employing a MINRES iterative solver yielded a speedup of 2×2$\times$2 × in floating point operations. These results highlight the significant potential for energy-aware scientific computing. The potential for future work lies in generalising this approach to other uncertainty quantification frameworks, such as multilevel Markov chain Monte Carlo or the multilevel stochastic collocation method and exploring its broader applications in scientific computing. The efficient parallel implementation of low-precision solvers within multilevel sampling methods is of great interest and promises to bring further speedup.